OpenStudy (anonymous):

x2 = 14x – 49 solve showing with work

OpenStudy (anonymous):

\[First convert \to ax^{2}+bx+c=0\] x^2-14x+49 = 0 a=1 b=-14 c=49 Plug into quadratic equation. \[x = (-b \pm \sqrt{b ^{2}-4ac})/2a\] x will have two answers

OpenStudy (vishweshshrimali5):

x^2 - 49 x + 49 = 0 => (x-7)^2 = 0 => x = 7 is the only sol.